19. Are you smarter than your older brother? In a study of birth order and intelligence, IQ tests were given to 18...

19.

Are you smarter than your older brother? In a study of birth order and intelligence, IQ tests were given to 18- and 19-year-old men to estimate the size of the difference, if any, between the mean IQs of firstborn sons and of secondborn sons. The following data for 10 firstborn sons and 10 secondborn sons are consistent with the means and standard deviations reported in the article. It is reasonable to assume that the samples come from populations that are approximately normal.

Firstborn

104

82

102

96

129

89

114

107

89

103

Secondborn

103

103

91

113

102

103

92

90

114

113

a.

Construct a 95% confidence interval for the difference in mean IQ between firstborn and secondborn sons.

b.

Based on the confidence interval, is it reasonable to believe that firstborn sons and secondborn sons may have the same mean IQ?
Answer #1
Values ( X ) Σ ( Xi- X̅ )2 Values ( Y ) Σ ( Yi- Y̅ )2
104 6.25 103 0.36
82 380.25 103 0.36
102 0.25 91 129.96
96 30.25 113 112.36
129 756.25 102 0.16
89 156.25 103 0.36
114 156.25 92 108.16
107 30.25 90 153.76
89 156.25 114 134.56
103 2.25 113 112.36
Total 1015 1674.5 1024 752.4

Mean X̅ = Σ Xi / n
X̅ = 1015 / 10 = 101.5
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 1674.5 / 10 -1 ) = 13.6402

Mean Y̅ = ΣYi / n
Y̅ = 1024 / 10 = 102.4
Sample Standard deviation SY = √ ( (Yi - Y̅ )2 / n - 1 )
SY = √ ( 752.4 / 10 -1) = 9.1433

Part a)

Confidence interval :-
(#1 – X2) ta/2,DF (S?/n1) + (S}/n2)
t(α/2, DF) = t(0.05 /2, 15 ) = 2.131

DF = ((S/n1+ s/n2))/((S/nl)/n1 - 1) + (S/n2)/n2 - 1))
DF = ((13.6402/10+9.1433/10))/((13.64022/10) /10-1)+(9.1433/10) /10-
DF = 15

(101.5 – 102.4) – t0.05/2,15 (13.64022/10) + (9.14332/10)
Lower Limit = (101.5 – 102.4) – t005/2,15 (13.64022/10) + (9.14332/10)
Lower Limit = -11.9659
Upper Limit = (101.5 – 102.4) + 0.05/2,15 (13.64022/10) + (9.14332/10)
Upper Limit = 10.1659
95% Confidence interval is ( -11.9659 , 10.1659 )

Part b)

Since the interval contain 0, we can conclude that  firstborn sons and secondborn sons may have the same mean IQ.